OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x − 3y = −3 2x − 6y = −6

OpenStudy (lynfran):

which method u prefer to used?

OpenStudy (anonymous):

Solve both equations for x. Then set the equations equal to each other and solve for y.

OpenStudy (lynfran):

ok using elimination method 5x-3y=-3...eq.1 2x-6y=-6...eq.2 we can multiply eq.2 by 2.5 to get 5x-15y=-15...eq..3 now we have 5x-3y=-3 5x-15y=-15 we can take away the two equations to solve for y...

OpenStudy (lynfran):

so we have 5x-5x=? , (-3y)-(-15y)=?? and (-3)-(-15)=??